OpenStudy (precal):

Find y' y=sinx + (4cosx)/x

OpenStudy (precal):

Do I do quotient rule for the 2nd part?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do u want me to show u?

OpenStudy (precal):

Let's see if I know how

OpenStudy (anonymous):

lol ok so u do it and then come back with the answer

OpenStudy (precal):

y'=cosx + x(-4sinx)-(4cosx)(1) y'=cosx-4sinx-4cosx y'=-3cosx - 4sinx

OpenStudy (precal):

did I do it right?

OpenStudy (anonymous):

no

OpenStudy (precal):

I forgot x^2 under it

OpenStudy (anonymous):

yes

OpenStudy (precal):

ok I will do it again

OpenStudy (anonymous):

but otherwise it is basically correct

OpenStudy (precal):

y'=cosx+ (-4sinx-4cosx)/x^2

OpenStudy (anonymous):

umm no you it shld be -4xsinx

OpenStudy (anonymous):

u included it the first time around

OpenStudy (precal):

you are so right y'=cosx+ (- 4xsinx-4cosx)/x^2

OpenStudy (anonymous):

ok ya this is fine

OpenStudy (anonymous):

you cld techinically factor out a -4

OpenStudy (anonymous):

but it isnt necessary. only if u want ur answer to look a lil neater

OpenStudy (precal):

what should my final answer look like? I am always interested in making it look neater.

OpenStudy (anonymous):